OpenStudy (anonymous):

A voltage V across a resistance R generates a current: I=V/R If a constant voltage of 9 puts across a resistance that is increasing at rate of .2 ohms/sec when the resistance is 5 ohms, at what rate if the current changing?

OpenStudy (kenljw):

I = V/R dI/dt = V (-1/R^2) dR/dt dI/dt = 9 (-1/5^2) .2 = - 1.8/25